"solve the following equation for x"
5/x-1 = 2/x+5

Answer by edjones(8007)

5/(x-1) = 2/(x+5) I assume that this is the problem.
5(x+5)=2(x-1) multiply each side by (x-1)(x+5) to eliminate fractions.
5x+25=2x-2
3x=-27
x=-9
.
Ed